A ball of mass m makes a head-on elastic collision with a second ball (at rest) and rebounds with a speed equal to one-fourth its original speed. What is the mass of the second ball?


1
Expert's answer
2021-01-22T11:41:40-0500
"p_i=p_f\\mv=-\frac{1}{4}mv+MV\\MV=\frac{5}{4}mv"

"0.5mv^2=0.5m(0.25)^2v^2+0.5MV^2\\
0.5mv^2=0.5m(0.25)^2v^2+0.5(1.25)^2\frac{m^2}{M}v^2\\
M=0.6m"